OpenStudy (anonymous):

How to answer "the non homo linear equation with constant coefficients" differential equation EASILY?

OpenStudy (anonymous):

99% of people on this site are in 8th grade math lol. Check out chegg for diffy Q What book are you using?

OpenStudy (anonymous):

@silverxx

OpenStudy (anonymous):

@perl Do you know Diffy Q?

OpenStudy (anonymous):

@lilshane

OpenStudy (lilshane):

well what do you think

OpenStudy (lilshane):

@jordan123321

OpenStudy (jordan123321):

yeahh what do you think

OpenStudy (perl):

Post a question and i will look at it :)

OpenStudy (lilshane):

yeah i kinda need a question

OpenStudy (lilshane):

@perl can you think about making me a ambassor

OpenStudy (lilshane):

r u still there

OpenStudy (lilshane):

The first special case of first order differential equations that we will look is the linear first order differential equation. In this case, unlike most of the first order cases that we will look at, we can actually derive a formula for the general solution. The general solution is derived below. However, I would suggest that you do not memorize the formula itself. Instead of memorizing the formula you should memorize and understand the process that I'm going to use to derive the formula. Most problems are actually easier to work by using the process instead of using the formula.

OpenStudy (lilshane):

does tht help you at all

OpenStudy (lilshane):

@sierra!

OpenStudy (lilshane):

@perl

OpenStudy (jordan123321):

perl

OpenStudy (perl):

hi

OpenStudy (anonymous):

@perl Have you taken diffy Q before?

OpenStudy (anonymous):

@perl ???

OpenStudy (anonymous):

@perl are you in calc 1?

OpenStudy (anonymous):

(x^3D^3- 3x^2d^2 + 7xD - 8)y= 12x^2 ln^2 x

OpenStudy (anonymous):

@perl =)

OpenStudy (anonymous):

de by kells @Austin6i6

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

i think @perl is in calc 1

OpenStudy (anonymous):

yes yes i think so

OpenStudy (anonymous):

I only took up to calc 3 for computer science . Although i am debating on taking diffy q. Is it worth it?

OpenStudy (anonymous):

yes it is. diff eq would test your knowledge in integral and diff cal. Deq isnt hard just always remember the basic

OpenStudy (anonymous):

@silverxx Whats your major?

OpenStudy (anonymous):

\[\left(x^3D^3-3x^2D^2+7xD-8\right)y=12x^2\ln^2x\] Is this really your ODE? Because those coefficients are not constant.

OpenStudy (anonymous):

At any rate, we can still try to find a solution. This looks like an Euler-Cauchy equation to me, so we'll try to solve the associated homogeneous ODE using the substitution \(y=x^r\). I'll switch to prime notation for the derivatives: \[\begin{align*} 0&=x^3y'''-3x^2y''+7xy'-8y\\\\ &=x^3\left(r(r-1)(r-2)x^{r-3}\right)-3x^2\left(r(r-1)x^{r-2}\right)+7x\left(rx^{r-1}\right)-8x^r\\\\ &=\left(r(r-1)(r-2)x^{r}\right)-3\left(r(r-1)x^{r}\right)+7\left(rx^{r}\right)-8x^r\\\\ &=r(r-1)(r-2)-3r(r-1)+7r-8\\\\ &=r^3-6r^2+12r-8\\\\ &=r^3-3(2)r^2+3(2^2)r-2^3\\\\ &=(r-2)^3 \end{align*}\] With \(r=2\) having multiplicity \(3\), we have the characteristic solution \[y_c=C_1x^2+C_2x^2\ln x+C_3x^2\ln^2x\]

OpenStudy (anonymous):

The nonhomogeneous solution might take some work... On a whim, I would attempt to find that solution via reduction of order.

OpenStudy (anonymous):

yes @SithsAndGiggles reduction of order would be better

OpenStudy (anonymous):

i think you missed something.. ln^5 x /5 should be there after c3.. Answer's on the book

OpenStudy (anonymous):

Right, I only found the homogeneous solution, but I'm still not sure how to obtain the nonhomogeneous one. I get the feeling we could use undetermined coefficients after a substitution, but that seems dicey.

OpenStudy (anonymous):

Reduction of order: Let's try to find a solution of the form \(y(x)=z(x)\,x^2\ln^2x\). \[\begin{cases}\begin{align*} y'&=(x^2\ln^2x)z'+(2x\ln^2x+2x\ln x)z\\\\ y''&=(x^2\ln^2x)z''+(4x\ln^2x+4x\ln x)z'+(2\ln^2x+6\ln x+2)z\\\\ y'''&=(x^2\ln^2x)z'''+(6x\ln^2x+6x\ln x)z''\\ &\quad\quad+(6\ln^2x+18\ln x+6)z'+\left(4\frac{\ln x}{x}+\frac{6}{x}\right)z \end{align*}\end{cases}\] Substitute into the ODE: \[\begin{align*} 12x^2\ln^2x&=x^3\bigg[(x^2\ln^2x)z'''+(6x\ln^2x+6x\ln x)z''\\ &\quad\quad\quad+(6\ln^2x+18\ln x+6)z'+\left(4\frac{\ln x}{x}+\frac{6}{x}\right)z\bigg]\\ &\quad-3x^2\bigg[(x^2\ln^2x)z''+(4x\ln^2x+4x\ln x)z'+(2\ln^2x+6\ln x+2)z\bigg]\\ &\quad+7x\bigg[(x^2\ln^2x)z'+(2x\ln^2x+2x\ln x)z\bigg]-8(x^2\ln^2x)z\\\\ 12x^2\ln^2x&=(x^5\ln^2x)z'''+(3x^4\ln^2x+6x^4\ln x)z''+(x^3\ln^2x+6x^3\ln x+6x^3)z' \end{align*}\] Substituting \(t=z'\), we get the second order linear ODE, \[12x^2\ln^2x=(x^5\ln^2x)t''+(3x^4\ln^2x+6x^4\ln x)t'+(x^3\ln^2x+6x^3\ln x+6x^3)t\] I'm not sure where to go from here... I'm considering giving undetermined coefficients a shot.